# MetaFaith: Faithful Natural Language Uncertainty Expression in LLMs

### v1

@@ Abstract
Gabrielle Kaili-May Liu1 Gal Yona2 Avi Caciularu2 Idan Szpektor2Tim G. J. Rudner3Arman Cohan1 1Yale University  2Google Research  3New York University {kaili.liu, arman.cohan}@yale.edu
@@ 1 Introduction
Understanding and improving the faithful calibration of LLMs is crucial to ensuring user trust and LLM reliability. Yet the influence of model, task, and prompt properties on faithful calibration remains poorly understood, with isolated studies of individual models (Yona et al., 2024; Ghafouri et al., 2024) overlooking systemic patterns and failure modes. To this end, we present the first systematic and comprehensive study of the faithful calibration problem in LLMs. While prior work (Ghafouri et al., 2024; Harsha Tanneru et al., 2024; Yona et al., 2024) asks if LLMs exhibit faithful calibration, we aim to go one step further and ask specifically when. We benchmark faithful calibration of LLMs through a comprehensive array of experiments spanning 16 models, 10 datasets, 6 content domains, and 5 uncertainty elicitation prompts. Examining the impact of various factors including model size, model post-training, task type, content domain, and prompt approach, we provide the most extensive evidence of faithful miscalibration of LLMs to date. We additionally provide insight into the impact of 12 advanced prompt strategies toward improving such calibration, finding approaches such as few-shot exemplars to be helpful but insufficient to reach substantial systematic improvement. Moreover, we show that leading factual calibration approaches prove largely unhelpful toward improving the faithfulness of LLM uncertainty expression, instead degrading alignment.
To address this critical challenge, we propose MetaFaith (Fig. 2), a systematic procedure for constructing calibration prompts that can robustly improve faithful calibration of any instruction-following LLM. Drawing inspiration from human metacognition, MetaFaith uses a carefully designed master prompt to guide a generator LLM to produce calibration prompts incorporating metacognitive strategies. These strategies enable models to self-reflect on their intrinsic confidence, communicate this internal state fluently, and embed uncertainty as a core part of their answers. By applying calibration prompts as system instructions, MetaFaith systematically modulates LLMs’ linguistically expressed confidence in a black-box fashion without requiring expensive training or access to model weights. We showcase the efficacy of MetaFaith through extensive experiments on 14 models and 10 datasets, finding that MetaFaith improves faithfulness by up to 61% and generalizes robustly across models, tasks, and domains. As we show, MetaFaith consistently improves over advanced, per-dataset prompt strategies, while being generalizable with use of a single prompt across all datasets. We further verify our results via human annotations, finding that MetaFaith enables models to achieve a win rate of 83% over a simple uncertainty elicitation baseline.
@@ 2 Related Work
Confidence Calibration of LLMs. Confidence calibration (Guo et al., 2017) is a fundamental aspect of building trustworthy AI systems (Desai and Durrett, 2020; Si et al., 2023). Existing methods primarily consider calibration from a factual perspective, aligning confidence with task accuracy (Kamath et al., 2020; Jiang et al., 2021; Geng et al., 2024; Huang et al., 2024; Xia et al., 2025). Such approaches can be classified into at least eight broad methodological divisions. [^2] Early work for pre-trained LMs (Xiao et al., 2022; Chen et al., 2023) investigated methods such as mixup (Park and Caragea, 2022), temperature scaling (Jiang et al., 2021), and label smoothing (Desai and Durrett, 2020). We do not discuss these further, instead focusing on more relevant recent works. Assuming access to internal model weights (“white-box” access), one popular class of approach aims to obtain estimates by examining probabilities assigned to individual tokens (Duan et al., 2024), probing internal representations (Azaria and Mitchell, 2023; Burns et al., 2024), computing token- or sentence-level entropy (Huang et al., 2025b), or adopting steering methods (Liu et al., 2024; Hong et al., 2025; Zhou et al., 2025c). Another line of work assumes only access to model outputs (i.e. “black-box” access). For example, semantic methods explore confidence estimation based on semantic consistency (Meister et al., 2022; Kuhn et al., 2023; Grewal et al., 2024; Nikitin et al., 2024), while sampling approaches assess variability across multiple outputs for a particular input, leveraging self-consistency or multi-stage assessment as a proxy measure of confidence (Kadavath et al., 2022; Manakul et al., 2023; Becker and Soatto, 2024; Chen and Mueller, 2024; Kaur et al., 2024; Xiong et al., 2024). Yet another direction targets calibration indirectly by learning auxiliary models to predict uncertainty or correctness (Shrivastava et al., 2023; Shen et al., 2024). Other techniques include test-time ensembling (Hou et al., 2024), use of prompt ensembles (Jiang et al., 2023), training with uncertainty-augmented data samples (Lin et al., 2022; Chaudhry et al., 2024; Zhang et al., 2024a), or self-reported probabilistic uncertainty (Tian et al., 2023; Yadkori et al., 2024; Yang et al., 2024a; Zhao et al., 2024). While such approaches investigate internal confidence of LLMs, they fail to consider the end-to-end nature of confidence calibration and the impact of linguistic assertiveness on perceived uncertainty (Ghafouri et al., 2024). In contrast, we aim to address the incorporation of uncertainty into model outputs, requiring significantly more expressivity and more closely resembling human uncertainty communication.
@@
Faithful Calibration of LLMs. A few recent works (Kumar et al., 2024; Ghafouri et al., 2024; Yona et al., 2024) explore the alignment between LLMs’ intrinsic and expressed uncertainty, but use of narrow experimental settings restricts the generalizability of their findings. Yona et al. (2024) proposes faithful response uncertainty as an example-level metric to reliably quantify faithful calibration, but their investigation is limited to proprietary LLMs and two short-form QA datasets. Ghafouri et al. (2024) finds the relationship between intrinsic confidence and linguistic assertiveness to be weak for GPT-4o, but they focus the misinformation domain and rely on a fine-tuned assertiveness judge model which may not generalize to other settings. Concurrently, Kumar et al. (2024) investigates faithful calibration of several GPT models and two small open-source LLMs but is limited to multiple-choice response formats and models linguistic confidence expression via categorical uncertainty phrases, which significantly undercuts expressivity. In comparison, we explore a significantly broader design space, considering a diverse array of uncertainty elicitation strategies, tasks, and content domains, as well as both proprietary and open-source models, spanning across several model families, sizes, and training procedures. Our results reveal persistent challenges across models and tasks, thus contributing a holistic and comprehensive understanding of faithful calibration.
@@ 3.2 Measuring Linguistic Decisiveness
To quantify linguistic decisiveness, we follow prior works (Ghafouri et al., 2024; Yona et al., 2024; Ji et al., 2025) and employ a LLM-as-a-Judge approach. Given a text input $Q$ and response $R$, we first instruct an evaluator LLM to extract assertions $A_{1},\ldots,A_{N}$ from $R$ using a carefully constructed few-shot prompt (§A.1, Fig. 5) (Yona et al., 2024). Thereafter, another few-shot prompt (§A.2, Fig. 6) is used to assess the decisiveness of each assertion and obtain a decisiveness score between 0 and 1. We use Gemini-2.0-Flash as the LLM judge for assertion extraction and decisiveness scoring, setting all inference hyperparameters to their default values in the Gemini Developer API. We validate our selection of judge model and the quality of our LLM-based scores by comparing against human annotations (further details in §3.4).
@@ 3.3 Measuring Intrinsic Uncertainty
3.4 Correlation with Human Judgment
Since our motivation is to improve the reliability and interpretability of LLM expressions of uncertainty in user-facing settings, we aim to quantify decisiveness in a way that aligns with humans perception. To this end, we investigated use of several different judge LLMs and prompt variants before finalizing our setup. We considered Gemini-1.5-Flash, Gemini-1.5-Pro, Gemini-2.0-Pro, Gemini-2.0-Flash, GPT-4o-Mini, GPT-3.5-Turbo, and GPT-4o as potential judges. [^6] Models such as Gemini 2.5 had not yet been released at the time of our experimentation. Preliminary experiments with large open-source models yielded poor results. We additionally varied the decisiveness prompt by adapting the judgment instructions and decisiveness scoring examples utilized by Yona et al. (2024) and Ghafouri et al. (2024). We studied the alignment of each combination of LLM judge and scoring prompt versus human perception through two experiments.
@@ 4.1 Experimental Setup
Models. Our experiments evaluate a total of 16 leading open- and closed-source models, varying in size, family, and post-training: GPT-4o-Mini (OpenAI et al., 2024), Gemini-2.0-Flash (Google Gemini Team, 2025), Qwen2.5-Instruct (1.5B, 7B, 72B) (Qwen et al., 2025), Llama3.1-Instruct (8B, 70B) (Grattafiori et al., 2024), Llama3.3-Instruct (70B), OLMo2-1124-Instruct (7B, 13B) (OLMo et al., 2025), Tulu3 (8B, 70B) (Lambert et al., 2025), Tulu3-8B-SFT, Tulu3-8B-DPO, and base models Qwen2.5-7B and Llama3.1-8B. Aside from Gemini-2.0-Flash, all models provide access to log-probabilities of output tokens. For all models we set the max output length to 250 tokens and temperature to 1.0. Responses for uncertainty estimation are obtained via beam search (beam size of 20).
Datasets. We select a suite of 10 datasets spanning diverse categories including knowledge-intensive QA, answerability, hallucination detection, math reasoning, scientific knowledge, computer science, social science, and commonsense reasoning: PopQA (Mallen et al., 2022), SelfAware (Yin et al., 2023), SimpleQA (Wei et al., 2024), MATH (Hendrycks et al., 2021b), UMWP (Sun et al., 2024), SciQ (Johannes Welbl, 2017), MMLU (Hendrycks et al., 2021a), HaluEval (Li et al., 2023), ARC-Challenge (Clark et al., 2018), and SuperGLUE (Wang et al., 2019). These datasets represent diverse difficulty levels and question types. We sample 1000 examples (Yang et al., 2024a; Yona et al., 2024) from the test split of each dataset to avoid potential dataset size bias. Additional dataset details are in §B.1.
@@ 4.2 What Influences Faithful Calibration?
best 0.72 0.71 0.72 0.00 0.69 0.70 0.62 0.55 0.63 0.71 0.61
GPT-4o-Mini none 0.50 0.53 0.51 0.00 0.51 0.51 0.50 0.50 0.44 0.51 0.45
best 0.57 0.59 0.60 0.10 0.53 0.52 0.54 0.53 0.56 0.67 0.52
@@
Instructing models to exhibit uncertainty where appropriate improves faithfulness, but specific prompt wording is unimportant. We observe that prompting models to express uncertainty boosts cMFG by up to 0.2, but the impact of prompt wording is mixed across models, with the best cMFG scores resulting from different prompts per model. Across prompts and datasets, models exhibit weak correlation between decisiveness and confidence (Table 2). Even with the best prompt per dataset LLMs failed to effectively hedge answers when unconfident or convey uncertainty when confident, suggesting that while prompting models to express uncertainty is a viable path to improve faithful calibration, obtaining systematic improvements is difficult. Additional analysis of the relative impact of each elicitation prompt can be seen in §E.1.

### v2

@@ Abstract
Gabrielle Kaili-May Liu1  Gal Yona2  Avi Caciularu2 Idan Szpektor2Tim G. J. Rudner3Arman Cohan1 1Yale University  2Google Research  3University of Toronto {kaili.liu, arman.cohan}@yale.edu
@@ 1 Introduction
Understanding and improving the faithful calibration of LLMs is crucial to ensuring user trust and LLM reliability. Yet the influence of model, task, and prompt properties on faithful calibration remains poorly understood, with isolated studies of individual models (Yona et al., 2024; Ghafouri et al., 2024) overlooking systemic patterns and failure modes. To this end, we present the first systematic and comprehensive study of the faithful calibration problem in LLMs. While prior work (Ghafouri et al., 2024; Harsha Tanneru et al., 2024; Yona et al., 2024) asks if LLMs exhibit faithful calibration, we aim to go one step further and ask specifically when. We benchmark faithful calibration of LLMs through a comprehensive array of experiments spanning 19 models, 10 datasets, 6 content domains, and 5 uncertainty elicitation prompts. Examining the impact of various factors including model size, model post-training, task type, content domain, and prompt approach, we provide the most extensive evidence of faithful miscalibration of LLMs to date. We additionally provide insight into the impact of 12 advanced prompt strategies toward improving such calibration, finding approaches such as few-shot exemplars to be helpful but insufficient to reach substantial systematic improvement. Moreover, we show that leading factual calibration approaches prove largely unhelpful toward improving the faithfulness of LLM uncertainty expression, instead degrading alignment.
To address this critical challenge, we propose MetaFaith (Fig. 2), a systematic procedure for constructing calibration prompts that can robustly improve faithful calibration of any instruction-following LLM. Drawing inspiration from human metacognition, MetaFaith uses a carefully designed master prompt to guide a generator LLM to produce calibration prompts incorporating metacognitive strategies. These strategies enable models to self-reflect on their intrinsic confidence, communicate this internal state fluently, and embed uncertainty as a core part of their answers. By applying calibration prompts as system instructions, MetaFaith systematically modulates LLMs’ linguistically expressed confidence in a black-box fashion without requiring expensive training or access to model weights. We showcase the efficacy of MetaFaith through extensive experiments on 19 models and 10 datasets, finding that MetaFaith improves faithfulness by up to 61% and generalizes robustly across models, tasks, and domains. As we show, MetaFaith consistently improves over advanced, per-dataset prompt strategies, while being generalizable with use of a single prompt across all datasets. We further verify our results via human annotations, finding that MetaFaith enables models to achieve a win rate of 83% over a simple uncertainty elicitation baseline.
@@ 2 Related Work
Confidence Calibration of LLMs. Confidence calibration (Guo et al., 2017) is a fundamental aspect of building trustworthy AI systems (Desai and Durrett, 2020; Si et al., 2023). Existing methods primarily consider calibration from a factual perspective, aligning confidence with task accuracy (Kamath et al., 2020; Jiang et al., 2021; Geng et al., 2024; Huang et al., 2024; Xia et al., 2025). Such approaches can be classified into at least eight broad methodological divisions. [^2] Early work for pre-trained LMs (Xiao et al., 2022; Chen et al., 2023) investigated methods such as mixup (Park and Caragea, 2022), temperature scaling (Jiang et al., 2021), and label smoothing (Desai and Durrett, 2020). We do not discuss these further, instead focusing on more relevant recent works. Assuming access to internal model weights (“white-box” access), one popular class of approach aims to obtain estimates by examining probabilities assigned to individual tokens (Duan et al., 2024), probing internal representations (Azaria and Mitchell, 2023; Burns et al., 2024), computing token- or sentence-level entropy (Huang et al., 2025b), or adopting steering methods (Liu et al., 2024; Hong et al., 2025; Zhou et al., 2025c). Another line of work assumes only access to model outputs (i.e. “black-box” access). For example, semantic methods explore confidence estimation based on semantic consistency (Meister et al., 2022; Kuhn et al., 2023; Grewal et al., 2024; Nikitin et al., 2024), while sampling approaches assess variability across multiple outputs for a particular input, leveraging self-consistency or multi-stage assessment as a proxy measure of confidence (Kadavath et al., 2022; Manakul et al., 2023; Becker and Soatto, 2024; Chen and Mueller, 2024; Kaur et al., 2024; Xiong et al., 2024). Yet another direction targets calibration indirectly by learning auxiliary models to predict uncertainty or correctness (Shrivastava et al., 2023; Shen et al., 2024). Other techniques include test-time ensembling (Hou et al., 2024), use of prompt ensembles (Jiang et al., 2023), training with uncertainty-augmented data samples (Lin et al., 2022; Chaudhry et al., 2024; Stengel-Eskin et al., 2024; Zhang et al., 2024a), or self-reported probabilistic uncertainty (Tian et al., 2023; Yadkori et al., 2024; Yang et al., 2024a; Zhao et al., 2024). Finally, more recent works have turned to cognition-inspired approaches to estimate and calibrate LLM confidence (Singh et al., 2024; Wen et al., 2024). While all of these methods are effective toward investigating internal confidence of LLMs, they fail to consider the end-to-end nature of confidence calibration and the impact of linguistic assertiveness on perceived uncertainty (Ghafouri et al., 2024). In contrast, we aim to address the incorporation of uncertainty into model outputs, requiring significantly more expressivity and more closely resembling human uncertainty communication.
@@
Faithful Calibration of LLMs. Faithfulness is well-studied in LLMs (Jacovi and Goldberg, 2020; Lyu et al., 2024; Chen et al., 2025) and refers to the accuracy with which an explanation represents a model’s underlying reasoning process. With regard to faithful confidence expression, a few recent works (Kumar et al., 2024; Ghafouri et al., 2024; Yona et al., 2024) explore the alignment between LLMs’ intrinsic and expressed uncertainty, but use of narrow experimental settings restricts the generalizability of their findings. Yona et al. (2024) proposes faithful response uncertainty as an example-level metric to reliably quantify faithful calibration, but their investigation is limited to proprietary LLMs and short-form QA. Ghafouri et al. (2024) finds the relationship between intrinsic confidence and linguistic assertiveness to be weak for GPT-4o, but their methodology focuses on misinformation tasks. Concurrently, Kumar et al. (2024) investigates faithful calibration of several GPT models and two small open-source LLMs but is limited to multiple-choice response formats and models linguistic confidence expression via categorical uncertainty phrases, which significantly undercuts expressivity. In comparison, we explore a significantly broader design space, considering a diverse array of uncertainty elicitation strategies, tasks, and content domains, as well as both proprietary and open-source models, spanning across several model families, sizes, and training procedures. Our results reveal persistent challenges across models and tasks, thus contributing a holistic and comprehensive understanding of faithful calibration.
@@ 3.2 Measuring Linguistic Decisiveness
To quantify linguistic decisiveness, we follow prior works (Ghafouri et al., 2024; Yona et al., 2024; Ji et al., 2025) and employ a LLM-as-a-Judge approach. Given a text input $Q$ and response $R$, we first instruct an evaluator LLM to extract assertions $A_{1},\ldots,A_{N}$ from $R$ using a carefully constructed few-shot prompt (§A.1, Fig. 5) (Yona et al., 2024). Thereafter, another few-shot prompt (§A.2, Fig. 6) is used to assess the decisiveness of each assertion and obtain a decisiveness score between 0 and 1. We use Gemini-2.0-Flash as the LLM judge for assertion extraction and decisiveness scoring, setting all inference hyperparameters to their default values in the Gemini Developer API. We validate the judgment paradigm and the quality of our LLM-based scores by comparing against human annotations (further details in §3.4).
@@ 3.3 Measuring Intrinsic Uncertainty
Hedge Word Human-Annotated Median (IQR) Mean Decisiveness (Ours) Mean Decisiveness (Yona et al., 2024)
“Almost No Chance" 0.02 (0.01, 0.05) 0.03 0.91
“Highly Unlikely" 0.05 (0.05, 0.10) 0.06 0.81
“Improbable" 0.10 (0.05, 0.22) 0.12 0.81
“Little Chance" 0.10 (0.05, 0.15) 0.14 0.81
“Chances are Slight" 0.10 (0.10, 0.20) 0.15 0.43
“Unlikely" 0.20 (0.10, 0.30) 0.20 0.86
“We Doubt" 0.20 (0.10, 0.30) 0.23 0.77
“Probably Not" 0.25 (0.15, 0.30) 0.33 0.74
“About Even" 0.50 (0.50, 0.50) 0.55 0.81
“Better than Even" 0.60 (0.55, 0.60) 0.64 0.72
“Likely" 0.70 (0.65, 0.75) 0.71 0.80
“Probably" 0.70 (0.60, 0.75) 0.68 0.84
“We Believe" 0.75 (0.65, 0.85) 0.75 0.93
“Very Good Chance" 0.80 (0.75, 0.90) 0.75 0.86
“Highly Likely" 0.90 (0.80, 0.95) 0.90 0.92
“Almost Certain" 0.95 (0.90, 0.98) 0.93 0.92
Table 1: Comparison of our mean decisiveness scores for common hedge words vs. the median and IQR of human perceptions of probability (Fagen-Ulmschneider, 2023), as well as vs. decisiveness scores obtained via the methodology of Yona et al. (2024). Decisiveness scores obtained via our paradigm show strong agreement with the human judgments, and moreso than those of Yona et al. (2024).
@@
3.4 Validating the Decisiveness Scores
Correlation with Human Judgment. Since our motivation is to improve the reliability and interpretability of LLM expressions of uncertainty in user-facing settings, we aim to quantify decisiveness in a way that aligns with humans perception. To this end, we investigated use of several different judge LLMs and prompt variants before finalizing our decisiveness scoring setup. We considered Gemini-1.5-Flash, Gemini-1.5-Pro, Gemini-2.0-Pro, Gemini-2.0-Flash, GPT-4o-Mini, GPT-3.5-Turbo, and GPT-4o as potential judges. [^6] Models such as Gemini 2.5 had not yet been released at the time of our experimentation. Preliminary experiments with large open-source models yielded poor results. We additionally varied the decisiveness prompt by adapting the judgment instructions and decisiveness scoring examples utilized by Yona et al. (2024) and Ghafouri et al. (2024). We studied the alignment of each combination of LLM judge and scoring prompt versus human perception through two experiments.
@@ 3.4 Validating the Decisiveness Scores
PopQA SelfAware SimpleQA
none basic MetaFaith none basic MetaFaith none basic MetaFaith
G2F 0.90 (±0.22) 0.87 (±0.27) 0.90 (±0.21) 0.94 (±0.14) 0.94 (±0.15) 0.95 (±0.14) 0.77 (±0.34) 0.82 (±0.28) 0.80 (±0.28)
G4oM 0.74 (±0.33) 0.74 (±0.34) 0.74 (±0.38) 0.90 (±0.20) 0.88 (±0.20) 0.84 (±0.24) 0.63 (±0.33) 0.64 (±0.36) 0.64 (±0.38)
Q2.5-1.5B 0.48 (±0.22) 0.45 (±0.23) 0.47 (±0.22) 0.55 (±0.23) 0.54 (±0.22) 0.55 (±0.23) 0.41 (±0.24) 0.34 (±0.22) 0.41 (±0.21)
Q2.5-7B 0.73 (±0.26) 0.70 (±0.30) 0.72 (±0.36) 0.79 (±0.20) 0.73 (±0.19) 0.72 (±0.26) 0.72 (±0.23) 0.67 (±0.25) 0.71 (±0.26)
L3.1-8B 0.49 (±0.25) 0.43 (±0.31) 0.45 (±0.23) 0.60 (±0.21) 0.63 (±0.22) 0.63 (±0.21) 0.53 (±0.23) 0.41 (±0.24) 0.43 (±0.22)
L3.1-70B 0.34 (±0.20) 0.36 (±0.22) 0.36 (±0.30) 0.54 (±0.22) 0.54 (±0.21) 0.56 (±0.20) 0.47 (±0.19) 0.40 (±0.22) 0.46 (±0.20)
Table 2: Robustness of the confidence scoring methodology across prompts and datasets for representative models.
@@
Alignment with Human Decisiveness Scores. To further validate the efficacy of our final decisiveness scoring paradigm, we present the results of a third experiment adapted from Yona et al. (2024). Using a similar setup as before, we randomly sample 320 model outputs (PopQA, basic prompt, 20 samples per model) and rewrite each answer to use a hedge expression from Fagen-Ulmschneider (2023). We then score the answers’ decisiveness using our scoring paradigm and that of Yona et al. (2024), and compute for each paradigm the mean decisiveness score issued for answers using each hedge word; these scores are compared against the distribution of human-perceived probabilities (Fagen-Ulmschneider, 2023) for each hedge word. Results are reported in Table 1. It can be seen that our scores are highly consistent with human-annotated judgments. While the approach used by Yona et al. (2024) does well on hedge words annotated with decisiveness of 0.5 and above, it yields poor results below this threshold, and rank-order is often not preserved. In contrast, our method is able to capture decisiveness in a human-aligned fashion across the whole range.
3.5 Robustness of Confidence Estimation
To validate our use of Gemini-2.0-Flash to obtain consistency judgments for confidence estimation, we follow the analysis by Yona et al. (2024) and compare the LLM judgments versus human judgments. We compute confidence scores for 160 randomly selected examples from PopQA across models (10 per model, responses elicited with the basic prompt) based on consistency judgments from Gemini-2.0-Flash versus author-assigned labels. We observe a high Spearman correlation of 0.98 between the scores resulting from each approach, slightly higher than the correlation reported by Yona et al. (2024).
A key factor in the robustness of sampling-based confidence estimates is to ensure estimates are not trivially influenced by the stability of sampled model responses under different prompt approaches. To this end, we show empirically that the distribution of confidence scores obtained via the sampling paradigm used in our experiments is not meaningfully influenced by prompts, suggesting the improved faithfulness is not coming from changes in quantified internal confidence but rather from adjustments to linguistic decisiveness.
Table 2 summarizes the mean and standard deviation of per-model per-dataset confidence scores for a representative sample of models [^7] We abbreviate model names in Table 2 as follows: G2F (Gemini-2.0-Flash), G4oM (GPT-4o-Mini), Q2.5-1.5B (Qwen2.5-1.5B-Instruct), Q2.5-7B (Qwen2.5-7B-Instruct), L3.1-8B (Llama3.1-8B-Instruct), L3.1-70B (Llama3.1-70B-Instruct). and datasets, across the uncalibrated (none), simple uncertainty prompt (basic), and MetaFaith prompt settings. We observe that confidence levels are generally stable across all settings, indicating robustness to prompt approach and task domain, the key variables in our experiments. These results are in line with existing work showing sampled estimates are reliable across domains and models Kuhn et al. (2023); Manakul et al. (2023); Rivera et al. (2024); Tian et al. (2024). Moreover, the cMFG metric for faithfulness is designed Yona et al. (2024) to help limit the effect of the confidence distribution.
@@ 4.1 Experimental Setup
Models. Our experiments evaluate a total of 19 leading open- and closed-source models, varying in size, family, and post-training: GPT-5(-Mini) (OpenAI et al., 2024), Gemini-2.5-Flash (Google Gemini Team, 2025), Qwen2.5-Instruct (1.5B, 7B, 72B) (Qwen et al., 2025), Llama3.1-Instruct (8B, 70B) (Grattafiori et al., 2024), Llama3.3-Instruct (70B), OLMo2-1124-Instruct (7B, 13B) (OLMo et al., 2025), Tulu3 (8B, 70B) (Lambert et al., 2025), Tulu3-8B-SFT, Tulu3-8B-DPO, and base models Qwen2.5-7B and Llama3.1-8B. Results for GPT-4o-Mini and Gemini-2.0-Flash are additionally provided in §E.2. All non-Gemini models provide access to log-probabilities of output tokens. For all models we set the max output length to 250 tokens and temperature to 1.0. Responses for uncertainty estimation are obtained via beam search (beam size of 20).
Datasets. We select a suite of 10 datasets spanning diverse categories including knowledge-intensive QA, answerability, hallucination detection, math reasoning, scientific knowledge, computer science, social science, and commonsense reasoning: PopQA (Mallen et al., 2022), SelfAware (Yin et al., 2023), SimpleQA (Wei et al., 2024), MATH (Hendrycks et al., 2021b), UMWP (Sun et al., 2024), SciQ (Johannes Welbl, 2017), MMLU (Hendrycks et al., 2021a), HaluEval (Li et al., 2023), ARC-Challenge (Clark et al., 2018), and SuperGLUE (Wang et al., 2019). While we choose tasks representing a diverse difficulty levels, since faithful calibration is precisely important in difficult task settings (Kim et al., 2024), our focus leans toward more challenging datasets to ensure faithful responses are expected to require expressing uncertainty. We sample 1000 examples (Yang et al., 2024a; Yona et al., 2024) from the test split of each dataset to avoid potential dataset size bias. Additional dataset details are in §B.1.
@@
GPT-5 none 0.51 0.52 0.51 0.37 0.46 0.36 0.51 0.51 0.36 0.49 0.46
best 0.70 0.69 0.72 0.68 0.60 0.63 0.60 0.59 0.53 0.67 0.64
GPT-5-Mini none 0.51 0.51 0.50 0.46 0.51 0.51 0.39 0.39 0.40 0.46 0.46
best 0.71 0.65 0.62 0.6 0.65 0.54 0.58 0.39 0.54 0.67 0.60
Gemini 2.5 Flash none 0.51 0.51 0.51 0.42 0.52 0.47 0.50 0.41 0.50 0.46 0.48
best 0.69 0.64 0.65 0.57 0.64 0.52 0.57 0.45 0.69 0.67 0.61
@@
4.2 What Influences Faithful Calibration?